OpenStudy (anonymous):

Solve for x: 5x + 12 = -3x + 60 A. -4 B. 6 C. -6 D. 8

OpenStudy (anonymous):

It was B 5(6) + 12 = 42 -3(6) + 60 = 42 so it was 6

jimthompson5910 (jim_thompson5910):

5x + 12 = -3x + 60 5x+3x = 60-12 8x = 48 x = 48/8 x = 6